A basketball player takes a shot a a regulation basket, 10 feet off the floor. He releases the ball with an initial upward velocity of 19 feet per second from a height of 7 feet off the floor. How long does it take for the ball to pass through the basket? The vertical motion model is: h=-16T^2+vT+s, where h is the height infeet as a function of time, T is the time in motion in seconds, v is the initial velocity in feet per second, and s is the height in feet.
Answer by stanbon(75887)

A basketball player takes a shot a a regulation basket, 10 feet off the floor. He releases the ball with an initial upward velocity of 19 feet per second from a height of 7 feet off the floor. How long does it take for the ball to pass through the basket? The vertical motion model is: h=-16T^2+vT+s, where h is the height infeet as a function of time, T is the time in motion in seconds, v is the initial velocity in feet per second, and s is the height in feet.
------------
h(t) = -16t^2 + 19t + 7
When the ball passes thru the hoop it is 10 ft. above the floor.
so, let h = 10 ft.
-------------
-16t^2+19t+7=10
-16t^2+19t-3=0
Use the Quadratic Formula:
t=[-19+-sqrt(19^2-4*-16*-3]/-32
t=0.1875 or t=2
-------
The 1st time is when the ball is on the way up.
The 2nd time is when the ball is coming down to 10 ft.
--------
2nd time is the answer you want.
=========
Cheers,
Stan H.